# Attention Layers Add Into Low-Dimensional Residual Subspaces

### v2

@@ 5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
Figure 4: The number of dead features (left) and the intrinsic dimension (right) of each activation in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024), which provide a complete set of SAEs trained on attention output, MLP output, and residual stream  [^5] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output.. We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4. This observation suggests that dead features may stem from the low-rank geometry of the activation space.
@@ 5.2 Active Subspace Initialization for Sparse Autoencoders
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ denote the input dimension, $h$ the hidden dimension of the SAE, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the singular value decomposition (SVD):
$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d},$
where $U\in\mathbb{R}^{n\times d}$ is the left singular vector matrix, $\Sigma\in\mathbb{R}^{d\times d}$ is the diagonal matrix of singular values, and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors.
We select the top $d_{\text{init}}$ right singular vectors to form the active subspace and then initialize the SAE weights directly in this subspace:
$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}},\quad W_{D}\in\text{span}(V_{\text{active}}),\quad W_{E}=W_{D}^{\top},$
where $W_{E}$ is the encoder weight matrix and $W_{D}$ is the decoder weight matrix. Intuitively, ASI aligns the initial SAE parameters with the active directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace. As $d_{\text{init}}$ decreases from the full space dimension [^6] Setting $D_{init}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization. within a certain range, the number of dead features in the SAE rapidly drops, with a corresponding improvement in Delta LM loss [^7] This metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens. (Figure 5). Further ablation studies and implementation details are contained in Appendix B and Appendix I, respectively.
Figure 5: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions for activations with a full space dimension of 4096.
Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than TopK and AuxK. A similar trend is observed in its impact on Delta LM Loss. Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
@@
The technique maintains applicability to diverse architectural variants and activation functions, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4 and Appendix H.
@@ 5.3 Scaling Laws
To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix G.5).
(a) Loss vs. Number of Parameters.
(b) Number of Dead vs. Number of Parameters.
(c) Loss vs. Number of Alive Parameters.
Figure 7: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from Llama-3.1-8B. (A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK. Active Subspace Init with SparseAdam achieves the best at large scale. (B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales. Enhanced with SparseAdam, dead features can be reduced to less than 1%. (C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency. Details in Section 5.3.
@@
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix C.
5.4 Generalize to Sparse Replacement Models
Recent work by He et al. (2025) reports that Lorsa, a sparse replacement model for attention layers, exhibits a high proportion of dead parameters. We hypothesize that the low-rank structure of attention outputs contributes significantly to this phenomenon.
To evaluate this, we apply Active Subspace Initialization to Lorsa. This modification reduces the proportion of dead parameters significantly under identical hyperparameter settings (Figure 8), while also improving the reconstruction. Specifically, we initialize the decoder (corresponding to $W^{O}$ in Lorsa) within the active subspace of each original MHSA head, while the encoder (corresponding to $W^{V}$ in Lorsa) is aligned with the corresponding input-side directions of the decoder. Further implementation details are provided in Appendix I.
Figure 8: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
6 Discussion
When to Use Active Subspace Initialization
For activation sites that do not exhibit clear low-rank structures (e.g., residual stream), our method yields only limited gains (Appendix B.2). The singular value spectrum provides a more reliable indicator: when the spectrum decays rapidly in its tail, applying ASI is likely to be effective.
Causality between Low-Rank Structure and Dead Features
We observed a strong correlation between low-rank structure and the emergence of dead features (Section 5), but the causal mechanism remains unclear. It may be tied to optimization dynamics or feature competition, yet a precise explanation is lacking. A deeper understanding of this relationship could also explain why ASI is less effective under high-$\ell_{0}$ warm-up regimes (Appendix H). We leave this important line of inquiry to future work.
@@ 7 Conclusion
Appendix A Declaration of LLM Usage
We acknowledge the use of large language models (LLMs) to assist in the preparation of this manuscript. Specifically, ChatGPT and Claude were utilized during the writing and coding process for the exclusive purpose of improving text clarity, grammar, overall fluency, and for generating boilerplate code structures. All ideas, theoretical developments, experimental designs, results, analyses, and scientific conclusions are entirely our own. The LLMs acted solely as assistive tools and were not involved in any aspect of the intellectual or scientific work.
Appendix B Ablation Study
B.1 Active Subspace Init vs Random Subspace Init
B.2 Apply on Near-Full-Rank Activation
B.3 Choice of the Initial Dictionary Size $D_{init}$
As shown in Figure 5, $D_{init}$ is a hyperparameter with a wide range of acceptable values (from 256 to 2048). We hypothesize that the performance degradation at very low values occurs because an excessively small dictionary fails to span the subspace containing critical information needed for effective reconstruction.
Appendix C Stale Momentum as Another Root Cause of Dead Features
Appendix D Activation Sources
@@
We study three large language models of different families–Llama-3.1-8B [^8] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^9] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^10] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture. This allows us to assess the robustness of spectral properties under varying model training configurations.
@@
To investigate how dataset diversity affects activation spectra, we select three datasets with varying linguistic and domain characteristics: (1) SlimPajama, an English corpus comprising web text, books, and other sources; (2) RedPajamaGithub, a large-scale code corpus; and (3) CCI3-Data, a Chinese dataset with broad domain coverage.
@@
Unless otherwise specified, activations are extracted from model intermediate layers. For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed). We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
Appendix E More Low-Rank Result Across Different Models and Datasets
(a) Activation spectra for many samples from SlimPajama fed into pythia-2.8b.
(b) Activation spectra for many samples from CCI3-Data fed into Qwen3-8B.
(c) Activation spectra for many samples from RedPajamaGithub fed into Qwen3-8B.
@@ Appendix E More Low-Rank Result Across Different Models and Datasets
We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 11. Models include pythia-2.8b [^11] https://huggingface.co/EleutherAI/pythia-2.8b. Datasets include RedPajamaGithub [^12] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^13] https://huggingface.co/datasets/BAAI/CCI3-Data
(a) Intrinsic dimension with threshold $0.5$.
(b) Intrinsic dimension with threshold $0.9$.
(c) Intrinsic dimension with threshold $0.999$.
(d) Intrinsic dimension with threshold $0.9999$.
Figure 12: Comparison of intrinsic dimensions across different variance thresholds.
Appendix F Different Choose of Variance Threshold for Intrinsic Dimension
We use 0.99 as the variance threshold in the main text. We show other threshold chose make no influence to the conclusion in Figure 12. Attention outputs show low-rank structure consistently.
Appendix G SAE Training Details
G.1 Collecting Activations
We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> and <eos> tokens.
G.2 Initialization
G.3 Optimization
We train SAEs using the Adam optimizers, both with $lr=2*10^{-4}$, $batchsize=32768$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The differences in scaling law experiments are described in the Appendix G.5.
G.4 Jumprelu SAEs
We trained JumpReLU SAEs (Rajamanoharan et al., 2024) under two distinct hyperparameter configurations: one maintaining a consistently low $\ell_{0}$ value throughout training, and another where $\ell_{0}$ is gradually decreased from a higher initial value. Unless otherwise specified, all JumpReLU SAEs were trained using the same settings as Conerly et al. (2025), which corresponds to the latter configuration. The key modifications for the former setting are as follows: (1) we initialized the encoder bias to zero instead of applying the heuristic that equalizes feature activation counts at initialization, and (2) we kept the sparsity coefficient fixed rather than employing a global warm-up schedule. As a result, the $\ell_{0}$ sparsity level started at a relatively low value early in training. This design is critical to our approach: we observed that if the model remains in a high-$\ell_{0}$ regime (e.g., on the order of $d_{\text{model}}/2$) for an extended period before sparsity increases, the feature directions tend to drift away from the active subspace during this phase, thereby diminishing the effectiveness of our method (Appendix H.1).
G.5 Fixed Hyperparameters in Scaling Law
@@
We empirically set $k=50$ for a reasonable sparsity in scaling laws.
@@
We empirically set the batch size to 40964096—below the critical batch size—which differs from the setting of $32768$ used in other experiments in this paper. This reduction is necessary because training SAEs with 1 million features requires substantial memory, posing significant computational challenges.
@@
Dimension of Subspace for SAE Initialization
We use 768 for all experiments.

### v3

@@ 5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
To study how low-rankness affects the interpretability of attention, we adopt the same framework and dataset as the original LlamaScope study (He et al., 2024) to evaluate their SAEs trained on attention outputs, MLP outputs, and the residual stream [^6] Another prominent set of open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output.. We find that the number of dead features is strongly related to effective rank, as shown in Figure 5. This observation suggests that dead features may stem from the low-rank geometry of the activation space. We also train SAEs using different SAE hyperparameters and further systematically verify that this phenomenon is prevalent in Appendix F.
(a) Loss vs. #Total Features.
(b) #Dead Features vs. #Total Features
(c) Loss vs. #Alive Features.
Figure 8: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from the middle layer of Llama-3.1-8B. (A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK. Active Subspace Init with SparseAdam achieves the best at large scale. (B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales. Enhanced with SparseAdam, dead features can be reduced to less than 1%. (C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency. Details in Section 5.3.
@@ 5.2 Active Subspace Initialization for Sparse Autoencoders
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ denote the input dimension, $h$ the hidden dimension of the SAE, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$ with singular value decomposition $\widetilde{A}=U\Sigma V^{\top}$ and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors, we select the top $d_{\text{init}}$ singular vectors to define the active subspace:
$\small V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}}.$
To initialize the SAE within this subspace, we first randomly initialize the decoder weights $W_{D}\in\mathbb{R}^{h\times d}$ and then project their first $d_{\text{init}}$ columns onto the active subspace:
$\small W_{D}\leftarrow W_{D}\,V_{\text{active}}\,V_{\text{active}}^{\top},\qquad W_{E}=W_{D}^{\top}.$
where $W_{E}$ is the encoder weight matrix and $W_{D}$ is the decoder weight matrix. Intuitively, ASI aligns the initial SAE parameters with the effective directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace. As $d_{\text{init}}$ decreases from the full space dimension [^7] Setting $d_{\text{init}}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization. within a certain range, the number of dead features in the Attention Output SAE rapidly drops, with a corresponding improvement in Mean Square Error (MSE) and Delta LM loss [^8] Following He et al. (2024), this metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens. (Figure 6). We refer readers to Appendix E for full SAE training details. Additional ablations, including the random subspace initialization baseline and the effect of activation rank, are reported in Appendix J. Pseudocode for Active Subspace Initialization (ASI) is provided in Appendix K.
@@
The technique maintains applicability to diverse architectural variants and activation functions, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4 and Appendix L.
We conduct a statistical significance test in Appendix H to demonstrate the statistical significance of our conclusions. We validate the effectiveness of ASI across different layers, models, and datasets in Appendix I. We further compare the features of TopK and ASI in Appendix N, analyzing both the degree of monosemanticity and the behavior of SAE features in the dead subspace, to ensure that ASI increases the number of alive features while preserving feature quality and maintaining the dictionary’s coverage and reconstruction performance in the dead subspace.
Figure 9: At a fixed number of Lorsa heads ($n=32768$), ASI (TopK Lorsa with Active Subspace Init) achieves a better reconstruction-sparsity trade-off than TopK (standard TopK Lorsa) and AuxK (TopK Lorsa with auxiliary loss). A similar trend is observed in its impact on Delta LM Loss.
@@ 5.3 Scaling Laws
To assess scaling, we evaluate our method as the number of SAE features grows from 1616K to 11M, keeping other hyperparameters fixed (see Appendix E).
@@
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the momentum terms and parameters corresponding to non-zero gradients, SparseAdam avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 8(a),  8(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix M.
5.4 Generalizing to Sparse Replacement Models
Recent work by He et al. (2025) shows that Lorsa—a sparse replacement for MHSA—has a substantial fraction of dead parameters. We hypothesize this is partly due to initialization: standard random initialization ignore the low-dimensional active subspace of attention outputs (Section 4).
Applying ASI to Lorsa.
To test whether our subspace-based approach extends beyond SAEs, we incorporate ASI (Section 5.2) into Lorsa training. When training Lorsa to approximate a target MHSA, we partition Lorsa heads into such groups, matching the number of attention heads in the MHSA. For each group, we initialize the encoder and decoder matrices directly within the corresponding input and output active subspaces of the MHSA head it corresponds. In addition, we initialize each group’s Q/K weights from the MHSA Q/K parameters. Pseudo code is provided in Appendix K
Results.
This initialization sharply reduces dead parameters under identical hyperparameters (Figure 9) while improving reconstruction quality. Further Lorsa training details are in Appendix O.
6 Discussion and Limitations
Low-rank attention outputs suggest a new direction for model improvement.
Modern large language models are predominantly built by stacking attention and MLP blocks with residual connections. Our finding that attention outputs show a low-rank structure suggests a potentially new avenue for improving model capacity and expressivity: future architectures may benefit from explicitly mitigating this rank collapse within attention modules. Notably, this phenomenon persists across models with and without grouped-query attention (GQA), and under both relative and absolute positional encodings (Section 4), indicating that it may reflect a more fundamental property of the attention mechanism itself rather than an artifact of specific design choices.
Causality between Low-Rank Structure and Dead Features.
We find a strong correlation between low-rank activations and the emergence of dead features (Section 5), but the underlying causal mechanism is unresolved. This effect may arise from optimization dynamics or feature competition, and we leave a rigorous explanation to future work.
When to Use Active Subspace Initialization.
Active Subspace Initialization is most beneficial when activations exhibit pronounced low-rank structure, such as attention outputs, residual streams of some narrow datasets, or certain specialized situations. For activation sites without clear low-rank behavior, the improvements are marginal (Appendix J.2).
@@ 7 Conclusion
Appendix A Activation Sources
@@
We study four large language models of different families–GPT-2999https://huggingface.co/openai-community/gpt2, Llama-3.1-8B [^10] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^11] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^12] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture. This allows us to assess the robustness of spectral properties under varying model training configurations.
@@
To investigate how dataset diversity affects activation spectra, we select two datasets with varying linguistic and domain characteristics: (1) SlimPajama [^13] https://huggingface.co/datasets/cerebras/SlimPajama-627B, an English corpus comprising web text, Github, Arxiv and other sources. Multi-Corpora in Figure 2 denotes the setting where all SlimPajama components are jointly used with random mixing. Github and Arxiv correspond to the code and scientific paper subsets of SlimPajama, respectively. (2) CCI3-Data [^14] https://huggingface.co/datasets/BAAI/CCI3-Data, a Chinese dataset with broad domain coverage, which is used in Appendix D as a supplement.
@@
We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
Appendix B Formal Definition of Fraction of Loss Recovered
Given the original language model cross-entropy loss is $loss_{\text{original}}$, the loss after ablating the activation at a specific position to zero is $loss_{\text{zero}}$, and the loss after replacing the original activation projected to the subspace spanned by first $n$ singular vectors is $loss_{\text{recovered}}$. Then, for these $n$ components, the fraction loss recovered is calculated as:
$\frac{loss_{\text{zero}}-loss_{\text{recovered}}}{loss_{\text{zero}}-loss_{\text{original}}}$
Appendix C Error analysis in Singular Value Decomposition
For the attention output of layer 15 of Llama-3.1-8B, we performed 5 times of singular value decompositions, using different 10 million tokens for each, and calculated the Coefficient of Variation (CV) for each singular value across these 5 runs. The maximum CV was only $4.9\times 10^{-3}$, and the mean and standard deviation of the effective rank computed from these 5 SVD results were 2523.165 and 0.404, respectively, with a CV of $1.5\times 10^{-4}$. These error experiments show that using 10 million tokens for singular value decomposition is sufficiently stable.
Appendix D More Singular Spectrum and Effective Rank Results
(a) Middle layer of pythia-2.8b; SlimPajama
Effective Dimensionality: Attention Output 1670;
MLP Output 2252; Residual Stream 2327
(b) Middle layer of Qwen3-8B; CCI3-Data
Effective Dimensionality: Attention Output 2356;
MLP Output 3558; Residual Stream 3140
(c) Middle layer of Qwen3-8B; SlimPajamaGithub
Effective Dimensionality: Attention Output 2410;
MLP Output 3495; Residual Stream 2000
(d) Middle layer of Qwen3-4B; SlimPajama
Effective Dimensionality: Attention Output 1122;
MLP Output 1485; Residual Stream 990
(e) Middle layer of gpt2; SlimPajama
Effective Dimensionality: Attention Output 515;
MLP Output 703; Residual Stream 656
(f) Middle layer of pythia-160m; SlimPajama
Effective Dimensionality: Attention Output 434;
MLP Output 594; Residual Stream 662
@@ Appendix D More Singular Spectrum and Effective Rank Results
D.1 Across Models and Datasets
We present relative singular values for some other model-dataset pairs in Figure 10. Models include pythia-160m [^15] https://huggingface.co/EleutherAI/pythia-160m, pythia-2.8b [^16] https://huggingface.co/EleutherAI/pythia-2.8b. Datasets include SlimPajamaGithub (subset of SlimPajama) and CCI3-Data.
D.2 Across Layers and Activation Positions
We present the effective rank for activations that are commonly used to train SAEs in Figure 11, including Concatenated Outputs of all Attention Heads (Z), attention output, the hidden activations of MLP (post activation fuction), MLP output, residual stream. All effective ranks are computed on SlimPajama.
(a) The number of the effective rank of each activation in Llama-3.1-8B
(b) The number of the effective rank of each activation in Gemma-2-9B
(c) The number of the effective rank of each activation in Qwen3-8B
Figure 11:
Appendix E SAE Training Details
We train SAEs as the following description.
E.1 Hyperparameters
@@
We empirically set $k=50$ for a reasonable sparsity following He et al. (2024), except the experiments for sweeping $k$.
Dictionary Size
We empirically set $n_{features}=32768$ which is $8\times d_{model}$, except the experiments for sweeping dictionary size (scaling law).
@@
We empirically set the batch size to $4096$.
Optimizer
We use the Adam and SparseAdam optimizer, both with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. Unless otherwise specified, Adam is used by default.
@@
We follow Gao et al. (2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512. We also sweep $\alpha$ and find the results are less sensitive to $\alpha$ in a reasonable interval.
Dimension of Subspace for SAE Initialization ($d_{init}$)
We use 768 for all experiments, except the experiments for sweeping $d_{init}$. We refer readers to Appendix J.3 for the reason.